How to solve using the substitution method
x^2+y=5
x-y+3=0

Answer by amarjeeth123(575)

The given equations are
x^2+y=5.........equation 1
x-y+3=0.........equation 2
From equation 2 we get y=x+3.....equation 3
Substituting in equation 1
x^2+x+3=5
x^2+x-2=0
x^2+2x-x-2=0
x(x+2)-1(x+2)=0
(x-1)(x+2)=0
x=1 or x=-2
y=4 or y=1
(1,4) and (-2,1) are the solutions.